OpenStudy (anonymous):

What is the simplified form of 24 y to the fifth power over 15 x to the eighth power divided by 8 y squared over 4 x to the fourth power ? 4 y cubed over 5 x to the fourth power 4 y to the fourth power over 5 x cubed 5 x to the fourth power over 4 y cubed 5 x cubed over 4 y to the fourth power

OpenStudy (anonymous):

can you help @StudyGurl14

OpenStudy (anonymous):

I will fan and medal

OpenStudy (studygurl14):

\[\frac{ \frac{ 24y^5 }{ 15x^8 } }{ \frac{ 8y^2 }{ 4x^4 } }\] like this?

OpenStudy (anonymous):